OpenStudy (anonymous):

Which of the following is not a perfect square trinomial? (Points : 1) 169 − 26y + y^2 81 + 18y + y^2 64 + 8y + y^2 25 + 10y + y^2

OpenStudy (lucaz):

you have to check them with Ax^2+Bx+C find two factors of A*C that add up to B here 169−26y+y^2 A*C=(1)*(169)=169, two factors of 169 that add up to -26? (-13)*(-13)=169 and -13+(-13)=-26, so the first option is a perfect square, (y-13)^2=169 − 26y + y^2 from (a+b)^2=ax^2+2ab+b^2, with neg. sign the 2ab term is also negative

OpenStudy (anonymous):

https://mathway.com/